OpenStudy (anonymous):

Choose the correct procedure used to graph the equation 15x - 5y = 30 using the intercepts method.

OpenStudy (anonymous):

you have to put it in your own words. here are the choices: A: Solve the equation for y to get a y-intercept of (0, -6) and a slope of 3. Plot the y-intercept, count the slope to get another point and draw the line that passes through the two points. B: Choose two different values; plug them into the equation to find the corresponding y-values. Plot the two points and draw the line that passes through the two points. C:Solve the equation for y to find an intercept of (-6, 0); solve the equation for x to find a y-intercept of (0, -2). Plot the two points and draw a line passing through the two points. D:Let x equal zero and solve for y to find a y-intercept of (0, -6). Let y equal zero and solve for x to find an x-intercept of (2, 0). Plot the two points and draw a line passing through the two points.

OpenStudy (anonymous):

first solve for the x-intercept (i.e. set y=0): 15x - 5(0) = 30 15x = 30 x = 2 the point is therefore (2, 0) then the y-intercept (set x=0): 15(0) - 5y = 30 -5y = 30 y = -6 the point is therefore (0, -6) this is effectively the same method as D.